# Flow-Anchored Consistency Models

### v1

@@ Abstract
Continuous-time Consistency Models (CMs) promise efficient few-step generation but face significant challenges with training instability. We argue this instability stems from a fundamental conflict: by training a network to learn only a shortcut across a probability flow, the model loses its grasp on the instantaneous velocity field that defines the flow. Our solution is to explicitly anchor the model in the underlying flow during training. We introduce the Flow-Anchored Consistency Model (FACM), a simple but effective training strategy that uses a Flow Matching (FM) task as an anchor for the primary CM shortcut objective. This Flow-Anchoring approach requires no architectural modifications and is broadly compatible with standard model architectures. By distilling a pre-trained LightningDiT model, our method achieves a state-of-the-art FID of 1.32 with two steps (NFE=2) and 1.76 with just one step (NFE=1) on ImageNet 256$\times$256, significantly outperforming previous methods. This provides a general and effective recipe for building high-performance, few-step generative models. Our code and pretrained models: https://github.com/ali-vilab/FACM.
Figure 1: Generated 1-step and 2-step samples on ImageNet 256$\times$256 by our FACM.
@@ 1 Introduction
As generative models scale to unprecedented sizes and applications demand real-time synthesis, the need for efficient, few-step samplers has become paramount. Consistency Models (CMs) (Song et al., 2023) have emerged as a promising paradigm for few-step generation. Early successful works were largely based on discrete-time formulations (Song et al., 2023; Song & Dhariwal, 2023; Geng et al., 2024), which are inherently prone to discretization errors. While their continuous-time counterparts can circumvent these errors, they have been historically hindered by severe training instability. Recent approaches, notably sCM (Lu & Song, 2024), have made significant strides in stabilizing continuous-time training through a combination of regularization techniques and architectural modifications. Concurrently, another line of research has aimed to stabilize training by reformulating the shortcut objective itself. For instance, IMM (Zhou et al., 2025) introduces additional self-consistency constraints between multiple timesteps, while MeanFlow (Geng et al., 2025) models the “average velocity” to arbitrary endpoints, implicitly learning about the underlying flow. Although these methods achieve stable few-step sampling, they do not explicitly decouple the task of modeling the instantaneous velocity field from the task of predicting the shortcut. We contend that this entanglement leads to sub-optimal performance, as the model must learn both the foundational flow and the shortcut simultaneously through a single, overly-coupled objective.
This paper addresses the root cause of instability in the continuous CM objective from a different perspective. We posit that the standard continuous CM objective, while powerful for learning a direct “shortcut” across a probability flow, is inherently unstable when trained in isolation. This is because it implicitly assumes the model has a robust understanding of the underlying flow, yet training exclusively on the shortcut objective can cause the model’s representation of this flow to degrade, leading to training collapse. Our key insight is that stability can be achieved by explicitly anchoring the model in the very flow it is shortcutting. The most direct way to achieve this Flow-Anchoring is to re-introduce the explicit training of the instantaneous velocity field that defines the flow. We propose that an objective based on Flow Matching (FM) (Lipman et al., 2022) can act as a crucial anchor, enabling the primary shortcut objective to be trained effectively. Based on this principle, we introduce the Flow-Anchored Consistency Model (FACM), which employs a simple yet effective training strategy combining two distinct objectives:
@@
  • Shortcut Objective that learns the efficient 1-step consistency mapping.
Our method is an architecturally-agnostic training strategy. We propose two effective implementations for this mixed-objective design, including an innovative Expanded Time Interval strategy that achieves optimal results without any architectural modifications. It supports both stable training from scratch and distillation from a pre-trained FM model, with the distillation pathway achieving superior performance at a significantly lower total training cost. Our approach sets new state-of-the-art FID scores of 1.76 (NFE=1) and 1.32 (NFE=2) on the class-conditional ImageNet 256$\times$256 benchmark, providing a general recipe for building high-performance, few-step generative models.
2 Preliminaries
Generative modeling aims to learn a probability flow ODE, $\mathrm{d}\bm{x}_{t}/\mathrm{d}t=\bm{v}_{\theta}(\bm{x}_{t},t)$, that transforms a prior distribution $p_{0}$ (e.g., $\mathcal{N}(0,I)$) to a data distribution $p_{1}$.
2.1 Flow Matching
Flow Matching (FM) (Liu et al., 2022; Albergo & Vanden-Eijnden, 2022; Lipman et al., 2022) learns the vector field $\bm{v}_{\theta}$ by regressing the network’s output against a conditional velocity field. This is achieved by defining a probability path between a noise sample $\bm{x}_{0}\sim p_{0}$ and a data sample $\bm{x}_{1}\sim p_{1}$. A simple and effective choice is OT-FM defined as $\bm{x}_{t}=(1-t)\bm{x}_{0}+t\bm{x}_{1}$, which has a constant conditional velocity of $\bm{x}_{1}-\bm{x}_{0}$. By training the model to predict this easily computable conditional velocity, it learns the marginal velocity field. This leads to the practical Flow Matching objective:
@@ 2.1 Flow Matching
2.2 Consistency Models
Consistency Models (CMs) (Song et al., 2023) are trained to map any point $\bm{x}_{t}$ on a given ODE trajectory directly to the trajectory’s endpoint in a single step. This is enforced through a self-consistency property: for any two points $\bm{x}_{t}$ and $\bm{x}_{t}^{\prime}$ on the same trajectory, their outputs must be identical, i.e., $f_{\theta}(\bm{x}_{t},t)=f_{\theta}(\bm{x}_{t}^{\prime},t^{\prime})$. This implies that for an ideal model, $f_{\theta}(\bm{x}_{t},t)$ should equal the trajectory’s endpoint $\bm{x}_{1}$ for all $t\in[0,1]$, which imposes a boundary condition $f_{\theta}(\bm{x}_{1},1)=\bm{x}_{1}$. CMs can be trained in both discrete and continuous time. Discrete-time CMs enforce consistency between two adjacent points with a finite distance, $f_{\theta}(\bm{x}_{t},t)\approx f_{\theta^{-}}(\bm{x}_{t+\Delta t},t+\Delta t)$, an approach that is inherently prone to discretization errors. Our work focuses on the continuous-time formulation. Here, the consistency property requires the total derivative of the shortcut function to be zero, $\frac{\mathrm{d}f_{\theta}(\bm{x}_{t},t)}{\mathrm{d}t}=0$. To satisfy the boundary condition, the function is parameterized as $f_{\theta}(\bm{x}_{t},t)=\bm{x}_{t}+(1-t)\bm{F}_{\theta}(\bm{x}_{t},t)$, which implies that the network $\bm{F}_{\theta}$ must satisfy:
@@ 2.2 Consistency Models
Here, $\bm{v}$ represents the conditional velocity $\bm{x}_{1}-\bm{x}_{0}$ from the underlying flow. In the distillation paradigm, this velocity is provided by a pre-trained FM teacher, which can also incorporate classifier-free guidance (CFG) (Ho & Salimans, 2022). This objective, which relies on the Jacobian-vector product (JVP) to compute the derivative term, is known to be unstable to train (Lu & Song, 2024).
@@ 3 Flow-Anchored Consistency Models (FACM)
This section first analyzes the core instability of continuous-time Consistency Models (CMs), identifying the “missing anchor” as the root cause. We then present our solution, the Flow-Anchored Consistency Model (FACM), detailing its mixed-objective training strategy.
3.1 Deconstructing CM Instability: The Missing Anchor
Our analysis reframes the challenge of training continuous-time Consistency Models. We argue that the instability is not an inherent flaw of the shortcut objective itself, but a consequence of training on it in isolation, which causes the model to lose its anchor in the flow’s underlying velocity field.
3.1.1 The Shortcut Target as an Average Velocity
@@
This term has a clear physical interpretation: it is the average velocity required to travel from point $\bm{x}_{t}$ to the endpoint $\bm{x}_{1}$ in the remaining time $1-t$. We denote this quantity as $\overline{\bm{v}}(\bm{x}_{t},t)$. Thus, the task of learning the 1-step shortcut is equivalent to training $\bm{F}_{\theta}$ to predict this average velocity.
@@
This identity is formally identical to the continuous-time CM learning objective (repeated from Eq. 2 for clarity):
This confirms that the CM objective directly forces the network $\bm{F}_{\theta}$ to learn the properties of an average velocity field, thus enabling the 1-step generation shortcut.
3.1.2 The Source of Instability: Losing the Flow Anchor
@@
Missing Instantaneous Velocity Field Supervision
The target $T$ explicitly depends on the ground-truth instantaneous velocity $\bm{v}$. The CM objective, however, only enforces a loss on the final prediction $\bm{F}_{\theta}$ (the average velocity). There is no explicit mechanism to ensure that the model’s learned dynamics remains faithful to the underlying instantaneous velocity field $\bm{v}$. The model is being asked to learn the integral of a function (average velocity) without being explicitly taught the function itself (instantaneous velocity).
Self-Referential Derivative Estimation
This lack of direct supervision on $\bm{v}$ makes the derivative term, $\frac{\mathrm{d}\bm{F}_{\theta^{-}}}{\mathrm{d}t}$, highly unstable. The total derivative, expanded via the chain rule, is:
@@
The network is required to estimate its own derivative. Even if the network is pre-trained (e.g., with Flow Matching) and initially provides a good approximation of the instantaneous velocity, the CM objective alone provides no continuous supervision to maintain this alignment. Without this anchor, the model’s output $\bm{F}_{\theta}$ can quickly drift, becoming meaningless or unstructured. In this state, the derivative term $\frac{\mathrm{d}\bm{F}_{\theta^{-}}}{\mathrm{d}t}$ becomes noisy and erratic. Using this noisy derivative to construct the training target creates a vicious cycle that rapidly amplifies errors and leads to training collapse.
These two issues are deeply intertwined, stemming from the same fundamental problem: the CM objective is ungrounded. It lacks a stable foundation in the very flow it is supposed to shortcut. The antidote is to re-introduce the explicit supervision of the instantaneous velocity field $\bm{v}$ via a Flow Matching objective. This provides a stable anchor for the model’s internal dynamics, ensuring that the model’s gradient field is well-behaved, which directly stabilizes the derivative term in the CM objective and allows the primary shortcut objective to be learned effectively. We term this principle Flow-Anchoring.
@@ 3.2.1 The FACM Objective: An Anchor and an Accelerator
Flow Matching (FM) Loss (The Anchor)
This loss component anchors the model by regressing its output towards the instantaneous velocity $\bm{v}$. The target $\bm{v}$ is constructed with a base velocity $\bm{v}_{\text{base}}$ and an optional classifier-free guidance (CFG) term:
@@
Consistency Model (CM) Loss (The Accelerator)
This component acts as an accelerator, training the model to learn the generative shortcut. We interpret the consistency condition (Eq. 6) as a fixed-point problem, $\bm{F}_{\theta}=T(\bm{F}_{\theta})$, where the operator is $T(\bm{F})\triangleq\bm{v}+(1-t)\frac{\mathrm{d}\bm{F}}{\mathrm{d}t}$. The training objective is designed to solve this problem stably and iteratively. First, we compute the consistency residual $\bm{g}$ of the stop-gradient model $\bm{F}_{\theta^{-}}$ ($\bm{F}_{\theta^{-}}=\text{sg}(\bm{F}_{\theta})$) :
$\bm{g}=\bm{F}_{\theta^{-}}(\bm{x}_{t},c_{\text{CM}})-T(\bm{F}_{\theta^{-}})=\bm{F}_{\theta^{-}}(\bm{x}_{t},c_{\text{CM}})-\left(\bm{v}+(1-t)\frac{\mathrm{d}\bm{F}_{\theta^{-}}(\bm{x}_{t},c_{\text{CM}})}{\mathrm{d}t}\right).$ (11)
This residual $\bm{g}$ is then clamped to the range $[-1,1]$ to prevent extreme gradients. A perturbed target is then formed as:
@@
$\bm{v}_{\text{tar}}=(1-\alpha(t))\bm{F}_{\theta^{-}}+\alpha(t)T(\bm{F}_{\theta^{-}}).$ (13)
This formulation provides a stable, interpolated learning target between the current model’s output and the ideal consistency target. The final CM loss component uses a norm L2 loss, $L_{\text{norm}}$ (detailed in Appendix A.2), and is modulated by weighting functions $\alpha(t)$ and $\beta(t)$:
@@
The combination of the interpolated target $\bm{v}_{\text{tar}}$ from the CM loss and the stabilizing flow anchor from the FM loss enables effective training. It is important to note that our specific choices for weighting and loss functions are designed to accelerate convergence, not as prerequisites for stability, which is guaranteed by the Flow-Anchoring principle.
@@ 3.2.2 Implementation of the Mixed Objective
A key design question is how to encode the distinct conditioning signals, $c_{\text{FM}}$ for the FM loss and $c_{\text{CM}}$ for the CM loss, that tell the model which velocity to predict. While this conditioning can include various information like class labels, for clarity in this section, we focus only on the time-based components. We explore two effective strategies for this (Figure 2):
Expanded Time Interval
We innovatively propose leveraging an expanded time domain to distinguish between the two tasks, a strategy that requires no architectural modifications. The primary CM task operates on the interval $t\in[0,1]$, using the time directly as the condition: $c_{\text{CM}}=t$. To perform the FM task at the same point $\bm{x}_{t}$ (defined by $t$), we signal this by mapping $t$ to the alternate interval $[1,2]$. This is done by setting the conditioning input to $c_{\text{FM}}=2-t$, which makes the two conditions decoupled, symmetric, and easily distinguishable. This mapping also ensures continuity at the boundary $t=1$, as the CM learning objective from Eq. 6 naturally converges to the FM objective’s target at the boundary:
@@
Auxiliary Condition with a Second Timestamp r
Alternatively, we can introduce a second time variable, $r$, to the model, making its full conditioning a tuple of $(t,r)$. We then define $c_{\text{CM}}=(t,1)$ and $c_{\text{FM}}=(t,t)$. This means the model signature is effectively $\bm{F}_{\theta}(\bm{x}_{t},t,r)$. When $r=1$, the model is trained on the CM task (predicting average velocity). When $r=t$, the model is trained on the FM task (predicting instantaneous velocity). We can provide this auxiliary condition to the model through a zero-initialized time embedder, which does not alter its original structure or initial output.
As shown in our ablations (Table 3), while both methods effectively stabilize training, the Expanded Time Interval strategy consistently yields the best performance. We attribute this to its use of highly distinct time domains ($[0,1]$ vs. $[1,2]$), which provide clearer, more separable conditioning signals for the two tasks compared to the subtler differences in the Auxiliary Time Condition (e.g., $(t,1)$ vs. $(t,t)$). The clarity of a signal appears to be more critical than the architectural change of adding a new input variable. Note that if the time schedule were $\bm{x}_{t}=t\bm{x}_{0}+(1-t)\bm{x}_{1}$, the conditions for the two strategies would be $t$ vs. $-t$ and $(t,0)$ vs. $(t,t)$, respectively.
@@
1:Online model $\bm{F}_{\theta}$, pretrained teacher $F_{\delta}$, metrics $d_{1},d_{2}$
@@
4:$\bm{x}_{t}\leftarrow(1-t)\bm{x}_{0}+t\bm{x}_{1}$
5:$\bm{v}\leftarrow\bm{x}_{1}-\bm{x}_{0}$triangleright For distillation, use $F_{\delta}(\bm{x}_{t},c_{\text{FM}})$ instead
6:$F_{\text{FM}}\leftarrow\bm{F}_{\theta}(\bm{x}_{t},c_{\text{FM}})$
7:$F_{\text{CM}},\nabla_{t}\bm{F}_{\theta}\leftarrow\texttt{JVP}(\bm{F}_{\theta},(\bm{x}_{t},c_{\text{CM}}),(\bm{v},1))$triangleright Simultaneous forward pass and JVP
8:$\overline{\bm{v}}\leftarrow\bm{v}+(1-t)\cdot\text{sg}(\nabla_{t}\bm{F}_{\theta})$
9:$\bm{v}_{\text{tar}}\leftarrow(1-\alpha(t))\cdot\text{sg}(F_{\text{CM}})+\alpha(t)\cdot\overline{\bm{v}}$triangleright Compute relaxation target
10:$\mathcal{L}_{\text{Total}}\leftarrow d_{1}(F_{\text{FM}},\bm{v})+d_{2}(F_{\text{CM}},\bm{v}_{\text{tar}})$
3.2.3 FACM Training Algorithm

### v2

# FACM: Flow-Anchored Consistency Models

@@ Abstract
[…]
Figure 1: FACM scales effectively to high-resolution text-to-image synthesis with a 14B parameter model (tops) and achieves state-of-the-art few-step generation on ImageNet $\times256$ (bottom).
@@ 1 Introduction
As generative models scale to unprecedented sizes and applications demand real-time synthesis, the need for efficient, few-step samplers has become paramount. Consistency Models (CMs) have emerged as a promising paradigm for few-step generation (song2023consistency). Early successful works were largely based on discrete-time formulations (song2023consistency; song2023improved; luo2023latent), which are inherently prone to discretization errors. While their continuous-time counterparts can circumvent these errors, they have been historically hindered by severe training instability. Recent approaches, notably sCM (lu2024simplifying), have made significant strides in stabilizing continuous-time training through a combination of regularization techniques and architectural modifications. Concurrently, Flow Mapping methods (geng2025meanflowsonestepgenerative; sabour2025align; wang2025transition) exemplify another line of research that has aimed to stabilize training. By reformulating the shortcut objective itself, these methods either model the “average velocity” to arbitrary endpoints, or introduce additional self-consistency constraints between multi-timesteps. Although these methods provide stable few-step sampling, they fail to address the root cause of instability. Their reliance on a single, over-coupled objective to learn the flow and shortcut simultaneously prevents explicit task decoupling and compromises perfect trajectory fidelity.
This paper addresses the root cause of instability in the continuous CM objective from a novel perspective. We posit that the standard continuous CM objective, while powerful for learning a direct “shortcut” across a probability flow, is inherently unstable when trained in isolation. This is because the approach implicitly assumes the model has a robust understanding of the underlying flow. However, training exclusively on the shortcut objective can induce catastrophic forgetting of this flow, leading to training collapse. Our key insight is that stability can be achieved by explicitly anchoring the model in the very flow it is shortcutting.
The most direct way to achieve this Flow-Anchoring is to re-introduce the explicit training of the instantaneous velocity field that defines the flow. We propose that an objective based on Flow Matching (FM) (lipman2022flow) can act as a crucial anchor, enabling the primary shortcut objective to be trained effectively. Based on this principle, we introduce the Flow-Anchored Consistency Model (FACM), which employs a simple yet effective training strategy combining two distinct objectives:
@@
  • Shortcut Objective that learns the efficient one-step consistency mapping.
[…]
2 Background
Diffusion and Flow Matching. Generative models aim to transform a prior distribution $p_{0}$ (e.g., $\mathcal{N}(0,I)$) to a data distribution $p_{1}$. A dominant approach is Diffusion Models (ho2020denoising; song2020denoising; karras2022elucidating), which learn to reverse a predefined noising process. Flow Matching (FM) (lipman2022flow; liu2022flow; albergo2022building; albergo2023stochastic; kingma2024understanding) offers a more direct framework to learn the probability flow ODE by regressing its output against a target velocity $\mathrm{d}\bm{x}_{t}/\mathrm{d}t=\bm{v}_{\theta}(\bm{x}_{t},t)$. A common approach uses the OT-FM path $\bm{x}_{t}=(1-t)\bm{x}_{0}+t\bm{x}_{1}$ between a noise sample $\bm{x}_{0}$ and a data sample $\bm{x}_{1}$, which has a constant conditional velocity of $\bm{x}_{1}-\bm{x}_{0}$. This leads to the practical FM objective:
@@ 2 Background
Consistency Models. Consistency Models (CMs) (song2023consistency) are trained to map any point $\bm{x}_{t}$ on an ODE trajectory directly to its endpoint $\bm{x}_{1}$ in a single evaluation. While early successful works were largely based on discrete-time formulations that are prone to discretization errors (song2023improved; geng2024consistency; luo2023latent; zheng2024trajectory), our work focuses on the continuous-time formulation. This approach requires the total derivative of the model’s output to be zero: $\frac{\mathrm{d}f_{\theta}(\bm{x}_{t},t)}{\mathrm{d}t}=0$. With the standard parameterization $f_{\theta}(\bm{x}_{t},t)=\bm{x}_{t}+(1-t)\bm{F}_{\theta}(\bm{x}_{t},t)$ and the boundary condition $f_{\theta}(\bm{x}_{1},1)=\bm{x}_{1}$, this implies the network $\bm{F}_{\theta}$ must satisfy:
@@
Here, $\bm{v}$ represents the conditional velocity $\bm{x}_{1}-\bm{x}_{0}$ from the underlying flow. In the distillation paradigm, this velocity is provided by a pre-trained FM teacher. This objective, relying on a Jacobian-vector product (JVP) for the derivative term, is notoriously unstable to train (lu2024simplifying). Recently, Flow Mapping methods (zhou2025inductive; geng2025meanflowsonestepgenerative; sabour2025align; wang2025transition; guo2025splitmeanflowintervalsplittingconsistency) have extended consistency models with a unified objective, but they do not address the root cause of instability and compromise perfect trajectory fidelity.
@@ 3 Flow-Anchored Consistency Models (FACM)
This section first analyzes the core instability of continuous-time Consistency Models (CMs), identifying the “missing anchor” as the root cause. We then present our solution, the Flow-Anchored Consistency Model (FACM), detailing its mixed-objective training strategy. Our analysis reframes the challenge of training continuous-time Consistency Models. We argue that the instability is not an inherent flaw of the shortcut objective itself, but a consequence of training on it in isolation, which causes the model to lose its anchor in the flow’s underlying velocity field.
3.1 Revisit the Shortcut Target of Consistency Models
@@
This term has a clear physical interpretation: it is the average velocity required to travel from point $\bm{x}_{t}$ to the endpoint $\bm{x}_{1}$ in the remaining time 1t1-t. We denote this quantity as $\overline{\bm{v}}(\bm{x}_{t},t)$. Thus, the task of learning the one-step shortcut is equivalent to training $\bm{F}_{\theta}$ to predict this average velocity.
@@
This identity is formally identical to the continuous-time CM learning objective (Eq. 2) and the Meanflow identity ($r\equiv 1$).
This confirms that the CM objective directly forces the network $\bm{F}_{\theta}$ to learn the properties of an average velocity field, thus enabling the one-step generation shortcut.
3.2 The Source of Instability: Losing the Flow Anchor
@@
Missing Instantaneous Velocity Field Supervision. The target $T$ explicitly depends on the instantaneous velocity $\bm{v}$. However, the CM objective only enforces a loss on the final prediction $\bm{F}_{\theta}$. There is no explicit mechanism to ensure that the model’s learned dynamics remain faithful to the underlying instantaneous velocity field $\bm{v}(\mathbf{x}_{t},t)$. The model is being asked to learn the integral of a function (average velocity) without being explicitly taught the function itself (instantaneous velocity).
Self-Referential Derivative Estimation. This lack of direct supervision on $\bm{v}$ makes the derivative term, $\frac{\mathrm{d}\bm{F}_{\theta^{-}}}{\mathrm{d}t}$, highly unstable. The total derivative, expanded via the chain rule, is:
@@
The network is optimized to estimate its own derivative to satisfy the consistency identity in Eq. 2. Ideally, this process should facilitate a smooth transition, evolving the model from predicting the instantaneous velocity field to an average velocity field that satisfies this identity. However, without a stable anchor in the underlying flow, the model’s output $\bm{F}_{\theta}$ quickly begins to drift. This drift has a critical consequence: the derivative term in the identity grows to dominate the ground-truth velocity $\bm{v}$, effectively diluting its supervisory signal. At this point, satisfying the identity no longer converges to the boundary condition. The training target thus becomes noisy and erratic, creating a vicious cycle that rapidly amplifies errors and ultimately leads to training collapse.
These two issues stem from the same fundamental problem: the CM objective is ungrounded. It lacks a stable foundation in the very flow it is supposed to shortcut. The antidote is to re-introduce the explicit supervision of the instantaneous velocity field $\bm{v}$ via a Flow Matching objective. This provides a stable anchor for the model’s internal dynamics, ensuring that the model’s gradient field is well-behaved, which directly stabilizes the derivative term in the CM objective and allows the primary shortcut objective to be learned effectively. We term this principle Flow-Anchoring.
@@ 3.3.1 The FACM Objective: An Anchor and an Accelerator
Flow Matching (FM) Loss (The Anchor). This loss component anchors the model by regressing its output towards the instantaneous velocity $\bm{v}$. The target $\bm{v}$ is constructed with a base velocity $\bm{v}_{\text{base}}$ and an optional classifier-free guidance (CFG) (ho2022classifier) term:
@@
Consistency Model (CM) Loss (The Accelerator). This component acts as an accelerator, training the model to learn the generative shortcut. We interpret the consistency condition (Eq. 2) as a fixed-point problem, $\bm{F}_{\theta}=\mathcal{T}(\bm{F}_{\theta})$, where the operator is $\mathcal{T}(\bm{F})\triangleq\bm{v}+(1-t)\frac{\mathrm{d}\bm{F}}{\mathrm{d}t}$. The training objective is designed to solve this problem stably and iteratively. First, we compute the consistency residual $\bm{g}$ of the stop-gradient model $\bm{F}_{\theta^{-}}$ ($\bm{F}_{\theta^{-}}=\text{sg}(\bm{F}_{\theta})$) :
$\bm{g}=\bm{F}_{\theta^{-}}(\bm{x}_{t},c_{\text{CM}})-\mathcal{T}(\bm{F}_{\theta^{-}})=\bm{F}_{\theta^{-}}(\bm{x}_{t},c_{\text{CM}})-\left(\bm{v}+(1-t)\frac{\mathrm{d}\bm{F}_{\theta^{-}}(\bm{x}_{t},c_{\text{CM}})}{\mathrm{d}t}\right).$ (10)
This residual $\bm{g}$ is then clamped to the range [1,1][-1,1] to prevent extreme gradients. A perturbed target is then formed as:
@@
$\bm{v}_{\text{tar}}=(1-\alpha(t))\bm{F}_{\theta^{-}}+\alpha(t)\mathcal{T}(\bm{F}_{\theta^{-}}).$ (12)
This formulation provides a stable, interpolated learning target between the current model’s output and the ideal consistency target. The final CM loss component uses a norm L2 loss, $L_{\text{norm}}$, and is modulated by weighting functions $\alpha(t)$ and $\beta(t)$ (detailed in Appendix A.3 and  A.4(c)):
@@
The combination of the interpolated target $\bm{v}_{\text{tar}}$ from the CM loss and the stabilizing flow anchor from the FM loss enables effective training. It is important to note that our specific choices for weighting and loss functions are designed to accelerate convergence, not as prerequisites for stability, which is already guaranteed by the Flow-Anchoring principle.
@@ 3.3.2 Implementation of the Mixed Objective
A key design question is how to encode the distinct conditioning signals, $c_{\text{FM}}$ for the FM loss and $c_{\text{CM}}$ for the CM loss, that tell the model which velocity to predict. While this conditioning can include various information like class labels, for clarity in this section, we focus only on the time-based components. We explore an effective strategy for this (Figure 2):
Expanded Time Interval. We innovatively propose leveraging an expanded time domain to distinguish between the two tasks, a strategy that requires no architectural modifications. The primary CM task operates on the interval $t\in[0,1]$, using the time directly as the condition: $c_{\text{CM}}=t$. To perform the FM task at the same point $\bm{x}_{t}$ (defined by $t$), we signal this by mapping $t$ to the alternate interval $[1,2]$. This is done by setting the conditioning input to $c_{\text{FM}}=2-t$, which makes the two conditions decoupled, symmetric, and easily distinguishable. This mapping also ensures continuity at the boundary $t=1$, as the CM learning objective from Eq. 2 naturally converges to the FM objective’s target at the boundary:
@@
Auxiliary Condition with a Second Timestamp. Alternatively, another intuitive approach is to introduce a second time variable, $r$, to the model, making its full conditioning a tuple of $(t,r)$. We then define $c_{\text{CM}}=(t,1)$ and $c_{\text{FM}}=(t,t)$. This means the model signature is effectively $\bm{F}_{\theta}(\bm{x}_{t},t,r)$. When $r=1$, the model is trained on the CM task (predicting average velocity from $t$ to 1). When $r=t$, the model is trained on the FM task (predicting instantaneous velocity at $t$, or from $t$ to $t$). We can provide this auxiliary condition $r$ to the model through a zero-initialized time embedder, which does not alter its original structure or initial output.
As shown in our ablations (Table 3), while both methods effectively stabilize training, the Expanded Time Interval strategy consistently yields the best performance. We attribute this to its use of highly distinct time domains ($[0,1]$ vs. $[1,2]$), which provide clearer, more separable conditioning signals for the two tasks compared to the subtler differences in the Auxiliary Time Condition (e.g., $(t,1)$ vs. $(t,t)$). For clarity, if $t=0$ represents the data distribution (i.e., $\bm{x}_{t}=t\bm{x}_{0}+(1-t)\bm{x}_{1}$), the conditions for the two strategies would be $t$ vs. t-t and $(t,0)$ vs. $(t,t)$, respectively.
@@
1:Online model $\bm{F}_{\theta}$, pretrained teacher $F_{\delta}$, metrics $\mathcal{L}_{\text{FM}},\mathcal{L}_{\text{CM}}$
@@
[…]
3.3.3 Training Algorithm and Scalable Chain-JVP Implementation